# Generative AI for Game Theory-based Mobile Networking

### v1

@@ Abstract
With the continuous advancement of network technology, various emerging complex networking optimization problems opened up a wide range of applications utilizating of game theory. However, since game theory is a mathematical framework, game theory-based solutions often require the experience and knowledge of human experts. Recently, the remarkable advantages exhibited by generative artificial intelligence (GAI) have gained widespread attention. In this article, we propose a novel GAI-enabled game theory solution that combines the powerful reasoning and generation capabilities of GAI to the design and optimization of mobile networking. Specifically, we first outline the game theory and key technologies of GAI, and then explore the advantages of combining GAI with game theory. Then, we briefly review the advantages and limitations of existing research and demonstrate the potential application values of GAI applied to game theory in mobile networking. Subsequently, we develop a game theory framework enabled by large language models (LLMs) to realize this combination, and demonstrate the effectiveness of the proposed framework through a case study in secured UAV networks. Finally, we provide several directions for future extensions.
@@ I Introduction
Game theory studies mathematical models of strategic interactions among interdependent rational participants (referred to as players), who make decisions based on their self-interests while their actions influence one another. Given its highly abstract representation of real-life situations, game theory has emerged as a standard analytical tool in various fields to model and comprehend complex systems involving multiple interacting agents. Employing game theory in the real world requires both constructing a suitable abstract game model (i.e., a binding from the cases to game-theoretic language) and solving the game (i.e., running an algorithm on the constructed model to extract the Nash equilibrium). Moreover, analyzing complex game scenarios with numerous players and intricate strategic interactions typically entails high computational complexity and communication costs among players. Additionally, game theory models often oversimplify real-world complexities for analytical convenience. However, this simplification may result in a model that fails to adequately capture all relevant factors in practical scenarios, thus affecting the accuracy and applicability of the model.
Solving the aforementioned issues can be challenging since it traditionally requires a deep understanding of game theory and solution algorithms, particularly for the newcomers or those with interdisciplinary backgrounds. Therefore, exploring the formulation of game-theoretic models from human language and obtaining a Nash equilibrium holds tremendous research potential, despite facing the following challenges.
  1. 1.
    Human language is not explicitly built into the foundations of game theory.
  2. 2.
    Human language often contains incomplete or vague information, making it difficult to precisely define all the necessary parameters of a game-theoretic model.
In particular, with the development of mobile networking, game theory plays an increasingly important role in network optimization, network configuration and management, as well as network security and privacy. For example, a game theory-based method [1] was proposed to optimize user allocation and transmission power to improve the performance of mobile edge computing networks. Therefore, exploring solutions to formulate game-theoretic models from human language and obtain their Nash equilibrium is also crucial to improve network performance and facilitate network management.
Fortunately, generative artificial intelligence (GAI) techniques, such as large language models (LLMs) combined with retrieval-augmented generation (RAG), offer a promising solution to address the aforementioned challenges. LLMs, such as GPT-4111https://openai.com/blog/chatgpt/, have demonstrated remarkable capabilities in understanding and generating human-like text. For example, the authors in [2] proposed a systematic simulation framework utilizing LLM agents for game theory research, and the simulation results demonstrated that the LLMs have the ability to imitate complex human strategic behaviors in socio-economic contexts. Furthermore, RAG models are able to enhance the reasoning capabilities of LLMs and provide more context-aware solutions by leveraging external knowledge sources, such as online encyclopedias or databases. Specifically, the advantages of using GAI in solving the challenges of formulating game models from human language and to obtain its Nash equilibrium in networking are as follows:
  • Learning from Data: Generative AI techniques allow for a data-driven approach to game model formulation. Specifically, LLMs can be trained on large-scale datasets, enabling them to learn from a wide range of game descriptions and strategies. This data-driven learning can help in capturing the patterns, common behaviors, and rationality assumptions exhibited by game players, their strategies and payoffs, enhancing the accuracy and realism of the formulated game models [3].
  • Incorporating External Knowledge: RAG models, which combine LLMs with retrieval-based methods, can leverage external knowledge sources to augment the generation process. This is particularly useful when dealing with incomplete or vague language descriptions. Furthermore, possessing the ability to learn from and access external data sources enhances adaptability. For example, we can leverage game-theoretic models derived from economics and effectively apply them to resource allocation in networking. By accessing relevant information from databases, RAG models can enhance the accuracy, context relevance, and credibility of game model formulation and solution. Moreover, they also improve the trustworthiness and interpretability of the formulated model.
  • Scalability and Computational Efficiency: LLMs and RAG models are designed with parallel processing capabilities, making them computationally efficient for handling large-scale game models. This is crucial when dealing with complex games involving numerous players and intricate interactions. Moreover, the efficiency of these models enables faster analysis and optimization, while reducing the computational burden associated with obtaining Nash equilibrium.
Motivated by these, this paper attempts to provide a forward-looking research that delves into the integration of GAI and game theory for networking optimization. To the best of the authors’ knowledge, the synergy between GAI and game theory for networking optimization has still been an open issue thus far. The contributions of this paper are summarized as follows:
@@
    We explore a few seminal applications of LLMs and game theory in networking from different perspectives, providing the guideline on how to integrated these three important technologies to solve practical problems.
@@
    We propose an LLM combined with RAG framework for networking optimization. Simulation results based on a real-world networking optimization case study validate the effectiveness of the proposed framework.
@@ II-B Generative AI Models
Generative Adversarial Networks (GANs): GANs are a type of neural network architecture with custom adversarial learning objectives, and they consists of two main parts that are the generator and discriminator. The goal of GANs is to train a generator network to generate samples that are similar to real data, and to train a discriminator network to distinguish between samples generated by the generator and real samples. Therefore, GANs have high performance in generating high-quality multi-modal content, which has led to a large number of applications such as 3D object generation, and image processing.
@@
Focus on LLM’s reasoning capabilities.
@@
Explore the use of LLM in strategic game experiments, specifically the ultimatum game and the prisoner’s dilemma.
Focus on the application of LLM in strategy game experiments.
@@
Focus on LLM’s collaborative capabilities.
@@
Explore the potential for LLM agents to spontaneously establish collaborative relationships in the absence of explicit instructions.
@@
Present one possible binding from dialogue to game theory as well as generalizations of existing equilibrium finding algorithms to this setting.
Focus on LLM-driven game generation.
II-C GAI for Game Theory
  • Computational Complexity: Obtaining Nash equilibrium for large games can incur significant computational costs. The search space for Nash equilibrium exponentially increases exponentially with the number of players and strategies [12]. Traditional AI techniques may struggle to efficiently explore the strategy space and converge to a Nash equilibrium in such cases.
LLMs combined with RAG provide a promising solution for addressing the abovementioned challenges faced by traditional AI. In the following, we elaborate on the advantages of LLMs combined with RAG in game scenarios from three perspectives, focusing on game scenario recognition, game theoretic model formulation, and Nash equilibrium search.
  • Nash Equilibrium Search: Obtaining the outcomes of game theoretic models, i.e., the Nash equilibrium, is often computationally complex. LLMs combined with RAG can extensively explore the strategy space in game theory and infer possible Nash equilibrium points. Moreover, LLM can enhance reasoning capabilities by learning existing methods, which makes obtaining Nash equilibrium more effective.
III LLMs-enabled Game Theory in Networking
In this section, we first provide a comprehensive review of the existing literature on the integration of LLMs and game theory, and analyze the contributions and limitations of the current works. Subsequently, we present some applications of LLMs-enabled game theory in networking and analyze the combination of LLM and game theory for multi-agent network design .
III-A Overview of Survey Papers
@@
Several studies have explored the integration of LLMs and game theory, as shown by Table I. However, these studies exhibit certain limitations. First, existing research mainly focuses on exploring the reasoning capabilities of LLMs in the context of game theory. However, the potential advantages of applying LLMs to game theory scenarios have not been well studied. Second, existing research is mainly based on typical game theory cases with pre-defined game theory models. However, in various application fields of game theory such as the networking, it is challenging to map actual game scenarios into appropriate game theory models. Given the potent natural language understanding and generation capabilities of LLMs, leveraging them to map game scenarios to game models is a great value topic, which is still lacking research. Finally, existing studies do not consider the combination of LLMs with RAG. However, RAG is crucial to improving the capabilities of LLMs. Therefore, the combination of LLMs and RAG needs further exploration. To this end, this paper proposes a framework that applies LLMs combined with RAG to game theory scenarios in networking to make up for the shortcomings of existing research.
@@
Figure 2: The LLM-enabled game theory framework. The framework is based on a layered architecture consisting of an input layer, an augmented layer, a decision layer, and an output layer. The input layer captures user input queries. The augmented layer employs RAG technology to enhance user queries. The decision layer utilizes a pluggable LLM to generate responses. The output layer returns the generated results to the user.
@@ III-B Potential Applications of LLMs-enabled Game Theory in Networking
III-C Combining LLM and Game Theory for Multi-Agent Network Design
The convergence of LLM, mobile networks, and multi-agent systems represents groundbreaking synergies that are leading to the emergence of multi-agent LLM network architectures, bringing huge potential for the design of future mobile networks [6]. This integration harnesses the power of collective intelligence and paves the way for self-managing networks. However, in a multi-agent LLM network, each LLM agent not only needs to pursue its individual objectives but also collaborate with other agents to achieve the collective goals of the entire network. This balance is crucial in designing efficient multi-agent LLM networks. In this context, game theory can be used to model and analyze the behavior of multi-agent LLM systems. This includes finding equilibria among agents, enabling them to serve individual objectives while effectively collaborating to achieve network-level goals. Furthermore, multi-agent reinforcement learning (RL) can further model the interactions among agents to learn optimal collaborative strategies and communication protocols among distributed LLM agents. In doing this, the communication costs among agents can be reduced.
IV LLMs-enabled Game Theory Framework Augmented by RAG for Mobile Network Optimization
Figure 3: The operation flow of RAG and LLM. In Part A, documents are loaded from the knowledge base, segmented into chunks, encoded into vectors, and stored in the vector database. Then, RAG retrieves the $K$ most relevant chunks to the query of user based on semantic relevance from the vector database. In Part B, the original query and the retrieved chunks are inputted together into LLM to generate the final response.
In the section, we first propose an LLM-enabled game theory framework supported by RAG for network optimization. Then, we conduct a case study on UAV secure communication optimization to demonstrate the effectiveness of the proposed framework.
@@ IV-A Motivations
Inspired by LLM’s outstanding capabilities, we propose an LLM-enabled game theory framework to address the abovementioned challenges. The proposed framework automates the mathematical formulation and game theory analysis of networking optimization problems through simple interactions with network designers, which effectively addresses the challenges of applying game theory and enhances the process of networking optimization. In addition, the framework can improve the accuracy of model formulation by incorporating relevant domain knowledge learned from a comprehensive knowledge base.
@@ IV-B Proposed Framework
  • Augmented Layer: The augmented layer is primarily implemented through the RAG technique, which is shown in detail in Part A of Fig. 3. The RAG can improve the quality, accuracy, and relevance of the generated answers of the system by providing relevant information related to the user requests [13]. The augmented layer contains a large-scale knowledge base to store rich material information, such as academic papers from IEEE Xplore. Specifically, RAG first acquires rich knowledge by loading the knowledge base, and then divides the knowledge into knowledge chunks that are suitable for embedded search. These knowledge chunks are then transformed into dense vector representations by using embedding techniques and stored in a vector database. Upon receiving the vector representation of the user request, RAG retrieves relevant information from the vector database and combines it with the user requests to form prompts, which are then sent to the decision-making layer for inference and decision-making execution.
  • Decision Layer: The decision layer adopts a plugin architecture to select the appropriate LLMs such as New bing, Bard, and GPT4, to formulate decisions. As shown in Part B of Fig. 3, LLM formulates a response based on the query and the selected chunks from the augmented layer. The decision layer mainly relies on the reasoning and decision-making capabilities of LLMs based on large-scale data training.
@@ IV-C Case Study: UAV Secure Communication Optimization
Figure 4: The experiment results of the wireless secure communication optimization case. The scenario module presents a system model of the considered scenario. The interaction module showcases the functionality of the proposed framework. The RAG augmentation module emphasizes the operational mechanism of RAG. The evaluation results module demonstrate the effectiveness of the proposed framework.
@@
Framework Configuration: In the framework configuration, we call the ChatGPT 4 model through the OpenAI API to implement the pluggable LLM module. Furthermore, the RAG module is built based on the LangChain.
@@
Framework Operation Flow: The LLM interactive module in Fig. 4 shows the optimization flow of the proposed framework to solve the UAV secure communication optimization problem. In traditional optimization process, a network designer needs to formulate a mathematical optimization problem, use an appropriate game theory model to analyze the problem, and design effective algorithms to solve it, which relies on rich professional knowledge. With the help of the proposed framework, the optimization process can be automatically implemented through three rounds of interaction between the network designer and the GAI agent. Therefore, the proposed framework can effectively solve the challenge faced by the traditional optimization process.
@@
Evaluation Results: The evaluation result module (a) of Fig. 4 shows the convergence of the algorithm generated by the proposed framework (i.e., GAI algorithm). From this figure, we can see that the GAI algorithm reaches a stable state, i.e., Nash equilibrium, after five iterations. The simulation results illustrate that the proposed framework can obtain a Nash equilibrium solution to the game theory problem. Moreover, the evaluation result module (b) of Fig. 4 shows the performance of the GAI algorithm with respect to different number of channels, where the expert algorithm indicates that the algorithm designed by network experts can be regarded as the optimal algorithm, and the EPR algorithm indicates that the average power allocation strategy is a benchmark algorithm. Correspondingly, we can observe that the GAI algorithm has the same performance as the expert algorithm. The evaluation results demonstrate the effectiveness of the proposed framework.

### v2

@@ Abstract
With the continuous advancement of network technology, various emerging complex networking optimization problems have created a wide range of applications utilizing game theory. However, since game theory is a mathematical framework, game theory-based solutions often rely heavily on the experience and knowledge of human experts. Recently, the remarkable advantages exhibited by generative artificial intelligence (GAI) have gained widespread attention. In this work, we propose a novel GAI-enabled game theory solution that combines the powerful reasoning and generation capabilities of GAI to the design and optimization of mobile networking. Specifically, we first outline the game theory and key technologies of GAI, and explore the advantages of combining GAI with game theory. Then, we review the contributions and limitations of existing research and demonstrate the potential application values of GAI applied to game theory in mobile networking. Subsequently, we develop a large language model (LLM)-enabled game theory framework to realize this combination, and demonstrate the effectiveness of the proposed framework through a case study in secured UAV networks. Finally, we provide several directions for future extensions.
@@ I Introduction
Game theory studies mathematical models of strategic interactions among interdependent rational participants (referred to as players), who make decisions based on their self-interests while their actions influence one another. Given its highly abstract representation of real-life situations, game theory has emerged as a standard analytical tool in various fields to model and comprehend complex systems involving multiple interacting agents. In particular, with the development of wireless technologies, game theory plays an increasingly important role in mobile networking optimization, such as network configuration and management, as well as network security and privacy protection [1]. For example, a game theory-based method [2] was proposed to optimize user allocation and transmission power to improve the performance of mobile edge computing networks. Employing game theory for mobile networking optimization requires both constructing a suitable abstract game model (i.e., a binding from the cases to game-theoretic models) and solving the game (i.e., running an algorithm on the constructed model to extract the Nash equilibrium). However, addressing the aforementioned issues can be challenging since it traditionally requires a deep understanding of game theory and solution algorithms, particularly for the newcomers or those with interdisciplinary backgrounds.
Fortunately, generative artificial intelligence (GAI) techniques, such as large language models (LLMs) combined with retrieval-augmented generation (RAG), offer a promising solution to address the aforementioned challenges. First, LLMs, such as ChatGPT-4, have demonstrated remarkable capabilities in natural language understanding and the generation of diverse content, with the potential to enable intelligent applications of game theory. Specifically, by offering natural language descriptions of network optimization problems, the LLMs have the potential to analyze these problems by using appropriate game theory models and generate effective solution algorithms, simplifying the use of game theory. Moreover, RAG technology can maintain a large-scale knowledge base that stores relevant knowledge on network optimization, game theory, and algorithms. By providing the LLMs with extensive specialized information, they can enhance the accuracy of content generation while overcoming the high dependency on specialized knowledge required for applying game theory.
Motivated by these, this work attempts to provide a forward-looking study to explore the use of LLMs and RAG technology to achieve the intelligent formulation and solution of game theory models by providing natural language descriptions of mobile networking optimization. To this end, we propose an LLM-enabled game theory framework, which implements the LLM module using the ChatGPT-4 model to understand human language, and enhances the generation ability of LLMs by building the RAG module using Langchain. To the best of our knowledge, this is the first work to systematically demonstrate the use of GAI technology to achieve the intelligence of game theory in mobile networking optimization. The contributions of this work are summarized as follows:
@@
    We explore a few seminal applications of LLMs and game theory in networking from different perspectives, providing the guideline on how to integrated these important technologies to solve practical problems.
@@
    We propose an LLM-enabled game theory framework for networking optimization. Simulation results based on a real-world networking optimization case study validate the effectiveness of the proposed framework.
@@ II-B Generative AI Models
Generative Adversarial Networks (GANs): GANs are a type of neural network architecture with custom adversarial learning objectives, which consist of a generator and a discriminator. The goal of GANs is to train a generator network to generate samples that are similar to real data, and to train a discriminator network to distinguish between samples generated by the generator and real samples. Therefore, GANs have high performance in generating high-quality multimodal content, which has led to a large number of applications such as 3D object generation, and image processing.
@@
II-C GAI for Game Theory
  • Computational Complexity: Obtaining Nash equilibrium for large games can incur significant computational costs. The search space for the Nash equilibrium increases exponentially with the number of players and strategies [6]. Traditional AI techniques may struggle to efficiently explore the strategy space and converge to a Nash equilibrium in such cases.
LLMs combined with RAG technology provide a promising solution for addressing the abovementioned challenges faced by traditional AI. In the following, we elaborate on the advantages of LLMs combined with RAG technology in game scenarios from three perspectives, focusing on game scenario recognition, game theoretic model formulation, and Nash equilibrium search.
  • Nash Equilibrium Search: Obtaining the outcomes of game theoretic models, i.e., the Nash equilibrium, often requires searching through large strategy spaces. Therefore, designing high-performance algorithms is crucial for reducing computational complexity. The combination of LLMs and RAG can generate high-performance algorithms by learning existing methods to obtain Nash equilibrium more efficiently.
III LLMs-enabled Game Theory in Networking
In this section, we first provide a comprehensive review of the existing literature on the integration of LLMs and game theory, and analyze the limitations of the current works. Subsequently, we present some applications of LLM-enabled game theory in networking and analyze the combination of LLMs and game theory for multi-agent network design.
III-A Overview of Survey Papers
Several studies have explored the integration of LLMs and game theory. Next, we will briefly review them and discuss their relevance to this work. Moreover, Table I provides a summary.
@@ III-A Overview of Survey Papers
Focus on LLMs’ reasoning capabilities.
@@
Explore the use of LLMs in strategic game experiments, specifically the ultimatum game and the prisoner’s dilemma.
Focus on the application of LLMs in strategy game experiments.
@@
Focus on LLMs’ collaborative capabilities.
@@
Explore the potential for LLM agents to spontaneously establish collaborative relationships in three competitive game scenarios.
@@
Present the integration of game-theoretic frameworks with LLMs, enhancing the strategic capabilities of LLMs in dialogue by modeling conversations as games.
Focus on the strategic capabilities of LLMs in dialogue.
Some works explored the reasoning capabilities of LLMs. For example, Fan et al. [9] evaluated the degree of rationality of LLMs in establishing clear desires, refining beliefs about uncertainty, and taking optimal actions in the context of game theory. However, the context of game theory is pre-established. Moreover, this work lacks an evaluation of more practical scenarios, such as network optimization. Duan et al. [8] analyzed the reasoning capabilities of LLMs in competitive environments through game-theoretic tasks that require pure logic and strategic reasoning to compete with opponents. However, this work utilized LLMs to generate actions for various game scenarios based on the predefined prompt templates, which can affect the outcomes due to the prompt sensitivity. Guo et al. [7] explored the use of LLMs in strategic game experiments, especially for the ultimatum game and the prisoner’s dilemma. However, these strategic game cases all have clear game theory models, and whether the LLMs can map real-world scenarios to these models remains to be explored.
There are also studies focused on evaluating the collaborative capabilities of LLMs. For example, Zou et al. [10] investigated the multi-agent LLMs from a game theoretic perspective, where the agents collaboratively solve tasks in the competitive environments. However, the proposed framework does not consider the issue of hallucinations in LLMs, where the generated content may be nonsensical or factually incorrect. In particular, model hallucinations may occur when the generative agents fail to capture the rapidly changing dynamics of wireless networks, leading to inaccurate representations of the wireless environment. Wu et al. [11] explored the potential for LLM agents to spontaneously establish collaborative relationships in three competitive game scenarios. However, this work only explores the potential of LLMs without incorporating the latest RAG technology. It is worth exploring how utilizing RAG to store extensive knowledge could enhance the ability of LLMs to establish collaborative relationships.
Moreover, Gemp et al. [12] investigated the integration of game-theoretic frameworks with LLMs, enhancing the strategic capabilities of LLMs in dialogue by modeling conversations as games. However, whether the strategic reasoning capabilities of LLMs in language understanding can be effectively applied to game-theoretic scenarios to enhance the practical applications of game theory remains underexplored.
@@
In general, the abovementioned studies have the following shortcomings compared to our work. First, the studies above mainly focus on exploring the reasoning capabilities of LLMs in the context of game theory. However, the potential advantages of applying LLMs to game theory scenarios have not been well studied. Second, the abovementioned studies are mainly based on classical game theory models with pre-defined game theory models. However, in various application fields of game theory such as the networking, it is challenging to map actual game scenarios in real world into appropriate game theory models. Given the potent natural language understanding and generation capabilities of LLMs, leveraging them to map game scenarios to game models is a great value topic, which is still lacking research. Finally, the abovementioned studies do not consider the combination of LLMs with RAG. However, RAG is crucial to improve the capabilities of LLMs. Therefore, the combination of LLMs and RAG needs further exploration. To this end, this paper proposes a framework that applies LLMs combined with RAG to game theory in networking scenarios to make up for the shortcomings of existing research.
@@
Figure 2: Part A shows the LLM-enabled game theory framework. The framework is based on a layered architecture consisting of an input layer, an augmented layer, a decision layer, and an output layer. The input layer captures user input queries. The augmented layer employs RAG technology to enhance user queries. The decision layer utilizes a pluggable LLM module to generate responses. The output layer returns the generated results to the user. Part B shows the flowchart of the proposed framework.
@@ III-B Potential Applications of LLMs-enabled Game Theory in Networking
III-C Combining LLMs and Game Theory for Multi-Agent Network Design
The convergence of LLMs, mobile networks, and multi-agent systems represents groundbreaking synergies that are leading to the emergence of multi-agent LLM network architectures, bringing huge potential for the design of future mobile networks [10]. This integration harnesses the power of collective intelligence and paves the way for self-managing networks. However, in a multi-agent LLM network, each LLM agent not only needs to pursue its individual objectives but also collaborate with other agents to achieve the collective goals of the entire network. This balance is crucial in designing efficient multi-agent LLM networks. In this context, game theory can be used to model and analyze the behavior of multi-agent LLM systems. This includes finding equilibria among agents, enabling them to serve individual objectives while effectively collaborating to achieve network-level goals. Furthermore, multi-agent reinforcement learning (RL) can further model the interactions among agents to learn optimal collaborative strategies and communication protocols among distributed LLM agents. In doing this, the communication costs among agents can be reduced.
IV LLM-enabled Game Theory Framework Augmented by RAG for Mobile Network Optimization
Figure 3: The operation flow of RAG models and LLMs. In Part A, documents are loaded from the knowledge base, segmented into chunks, encoded into vectors, and stored in the vector database. Then, RAG retrieves the $K$ most relevant chunks to the query of user based on semantic relevance from the vector database. In Part B, the original query and the retrieved chunks are inputted together into LLMs to generate the final response.
In the section, we first propose an LLM-enabled game theory framework for network optimization. Then, we conduct a case study on UAV secure communication optimization to demonstrate the effectiveness of the proposed framework.
@@ IV-A Motivations
Inspired by the outstanding capabilities of LLMs, we propose an LLM-enabled game theory framework that mainly consists of an LLM module and a RAG module. Specifically, the LLM module can automatically complete the mathematical formulation of network optimization problems, analyze the formulated optimization problems by using appropriate game theory models, and generate corresponding solutions through simple interactions with the user. Moreover, the RAG module contains a large-scale knowledge base that stores rich material data such as academic papers from IEEE Xplore, and it can improve the accuracy of LLM-generated responses by retrieving knowledge relevant to the user requests.
@@ IV-B Proposed Framework
  • Augmented Layer: The augmented layer is primarily implemented through the RAG technique, which is shown in detail in Part A of Fig. 3. The RAG can improve the quality, accuracy, and relevance of the generated answers of the system by providing relevant information related to the user requests [13]. Specifically, The application of RAG consists of two phases, i.e., the database preparation phase and the knowledge retrieval phase. 1) Database Preparation Phase: RAG contains a large knowledge base, which stores sufficient external information, such as academic papers from IEEE Explore, technical documents, and algorithms from users. The information above is first divided into smaller knowledge chunks through semantic chunking techniques to facilitate embedded retrieval. Subsequently, a semantic embedding model is used to encode these knowledge chunks into vector representations, which are stored in an efficient vector database. 2) Knowledge Retrieval Phase: When the user inputs a query, RAG uses the same embedding model for constructing the vector database to convert the user query into a vector representation, which is then used to retrieve the vector database. Subsequently, RAG ranks the retrieved results based on vector similarity and returns the most relevant knowledge chunks. These retrieved knowledge chunks along with the user query form a prompt, which is sent to the decision-making layer as a foundational input for reasoning and decision-making by the generative model.
  • Decision Layer: The decision layer adopts a plugin architecture, allowing it to select an appropriate LLM (such as New Bing, Bard, and GPT-4) to perform reasoning tasks according to the query. As shown in Part B of Fig. 3, the LLMs generates responses based on the query and the relevant knowledge blocks retrieved from the augmentation layer. This approach enables the LLMs not only to leverage its internal language understanding and generation capabilities, but also to dynamically access external knowledge, resulting in more precise responses.
@@ IV-C Case Study: UAV Secure Communication Optimization
Figure 4: The experiment results of the UAV secure communication optimization case. The scenario module presents a system model of the considered scenario. The interaction module showcases the functionality of the proposed framework. The RAG augmentation module emphasizes the operational mechanism of RAG. The evaluation results module demonstrate the effectiveness of the proposed framework.
@@
Framework Operation Flow: The LLMs interactive module in Fig. 4 shows the optimization flow of the proposed framework to solve the UAV secure communication optimization problem. In traditional optimization process, a network designer needs to formulate a mathematical optimization problem, use an appropriate game theory model to analyze the problem, and design effective algorithms to solve it, which relies on rich professional knowledge. With the help of the proposed framework, the optimization process can be automatically implemented through three rounds of interaction between the network designer and the GAI agent. Therefore, the proposed framework can effectively solve the challenge faced by the traditional optimization process.
Simulation Setup: To validate the effectiveness of the proposed framework, we conducted relevant simulation experiments. The specific simulation setup are described as follows.
  • Framework Setup: For the proposed framework, the pluggable LLM module is implemented by OpenAI APIs for calling the ChatGPT-4 model, and the RAG module is built based on LangChain.
  • Key Parameters: The size of chunks in the RAG is set to 2000, and the GAI agent retrieves 3 chunks to perform inference in each round. The total power of the friendly UAV jammer is set as $P_{\text{total}}=0.05\ \text{W}$, and the transmission power of the legitimate UAV users is set as $P_{s}=0.01\ \text{W}$. Moreover, the channel coefficients from the legitimate UAV users to the destination base station, from the legitimate UAV users to the eavesdropper and from the UAV jammer to the eavesdropper are set as $g_{s},g_{e},g_{e,i}\in[0,1]\times 10^{10}\ \text{W}$.
  • Evaluation Metrics: To validate the effectiveness of the proposed framework, we evaluated the following performance metrics. 1) Convergence: Convergence evaluates whether the algorithm generated by the proposed framework can achieve a Nash equilibrium solution. 2) System Sum Secrecy Rate: The system sum secrecy rate represents the optimization objective of the considered scenario, with a higher secrecy rate indicating better performance.
@@
Evaluation Results: The evaluation result module (a) of Fig. 4 shows the convergence of the algorithm generated by the proposed framework (i.e., GAI algorithm). From this figure, we can see that the GAI algorithm reaches a stable state, i.e., Nash equilibrium, after five iterations. The simulation results illustrate that the proposed framework can obtain a Nash equilibrium solution to the game theory problem.
Moreover, the evaluation result module (b) of Fig. 4 shows the performance of the GAI approach with respect to different number of channels, where the expert approach indicates that the algorithm designed by the network experts can be regarded as the optimal approach, the EPR approach indicates that the average power allocation strategy is a benchmark approach, and the multi-agent deep deterministic policy gradient (MADDPG)-based approach (i.e., the MADDPG-based approach) is a deep reinforcement learning method, which is used as a comparative approach. Correspondingly, we can observe that as the number of channels increases, the system sum secrecy rate shows an upward trend. This is because more legitimate UAV users participate in communication. Further, the GAI-based approach has approximately the same performance as the expert algorithm. This can be attributed to the fact that the proposed framework can generate accurate solution approaches for the considered scenario by learning from a large amount of relevant information.
The GAI-based approach outperforms the MADDPG-based approach and the EPR approach, and the reasons are as follows. First, due to the heterogeneous communication conditions of legitimate UAV users, the average power allocation strategy of EPR struggles to efficiently utilize the limited power resources. Second, the MADDPG-based approach requires agents to continuously interact and learn to discover the optimal strategy. This dynamic learning process makes it difficult for the agents to achieve effective strategy coordination. In conclusion, the evaluation results demonstrate the effectiveness of the proposed framework.
IV-D Discussion
In this section, we further discuss the proposed framework including the requirements and new challenges as well as the associated costs.
1) Requirements and Challenges: The proposed framework introduces the following requirements and novel challenges:
  • Accurate and Diverse Database: In the proposed framework, the LLMs need to be able to handle and understand diverse optimization problems and game theory models. Therefore, the accuracy and diversity of the database provided by RAG are crucial to ensure that the generated content is not only correct but also covers a wide range of application scenarios. However, obtaining accurate and diverse database is challenging.
  • Detailed and Precise Requests: The proposed framework necessitates interaction with the user to generate appropriate responses. Therefore, satisfactory generation outcomes require the user to provide detailed and precise requests. However, ensuring that the users can provide sufficiently detailed and precise requests is challenging.
  • Additional Parameter Tuning: The integration of LLMs and RAG models introduces additional parameters, such as the size of knowledge chunks and the number of retrieval blocks. These parameters not only influence the response time but also affect the accuracy and consistency of the LLMs. Therefore, achieving optimal performance requires careful parameter tuning tailored to specific optimization problems, which is also challenging.
2) Associated Costs: Implementing the proposed framework mainly involves the following costs:
  • Database Collection Costs: RAG necessitates maintaining a large-scale database to store extensive knowledge related to network optimization and game theory, including academic papers, technical documents, and algorithms, in order to enhance the accuracy of LLM-generated responses. Therefore, the construction of the database, including information collection and filtering, incurs significant costs.
  • Computational Costs: RAG extracts knowledge which is relevant to user requests from the large-scale database, and the LLMs generate responses to these requests. This process often requires substantial computational resources, leading to high computational costs.